OpenStudy (anonymous):

If 1/3 of a 12-foot board is sawed off, how much is left? 3 feet 6 feet 4 feet 8 feet

OpenStudy (mitu12):

1 X 12 3 X 1 12/3 = 4 4-12= 8

OpenStudy (anonymous):

how did the 3 turn to 4?

OpenStudy (mitu12):

12/3=4